OpenStudy (theraggedydoctor):

If \(\frac{ 1 }{ 2 }\) is \(\frac{ 3 }{ 4}\) of \(\frac{ 4 }{ 5 }\) of a certain number, what is that number?

OpenStudy (theraggedydoctor):

@TheSmartOne

OpenStudy (dmndlife24):

\[\frac{ 1 }{ 2 } = \frac{ 3 }{ 4 }\times \frac{ 4 }{ 5 } \times x\] Let x equal that certain number

TheSmartOne (thesmartone):

@dmndlife24 has it correct :)

TheSmartOne (thesmartone):

Just something that you should remember, the word 'of' in math means multiplication. 'is' means equal to.

OpenStudy (dmndlife24):

Simplify... \[\frac{ 1 }{ 2 } = \frac{ 3 }{ 5 }x\]

TheSmartOne (thesmartone):

and your certain number, you can make that any variable and solve for it :)

OpenStudy (dmndlife24):

@TheRaggedyDoctor Tell me what you get when you solve for "x" :)

OpenStudy (theraggedydoctor):

I'm not sure..

TheSmartOne (thesmartone):

Hint: Multiply both sides by the reciprocal of \(\sf\Large\frac{3}{5}\) The reciprocal of \(\sf\Large\frac{a}{b}\) is \(\sf\Large\frac{b}{a}\)

OpenStudy (theraggedydoctor):

\(\frac{ 6 }{ 5 }\)

OpenStudy (dmndlife24):

\[\frac{ 1 }{ 2 } \times \frac{ 5 }{ 3 } = x\]

OpenStudy (theraggedydoctor):

\(\frac{5}{6}\)

OpenStudy (dmndlife24):

Yup

OpenStudy (theraggedydoctor):

Oh okay. Thank you. c:

OpenStudy (dmndlife24):

Sure thing :)